# Murphy: Multi-Turn GRPO for Self Correcting Code Generation

### v1

@@ 4.1 Pruning Strategies in Murphy
Intra-Group Pruning (IntraP):In IntraP, pruning operates recursively: starting from the children of stage S1S-1, each pruning step is followed by reward propagation as described in Sec. 4, and the process continues backward until reaching the root. This ordering—pruning before reward reassignment—ensures that only informative trajectories are retained for credit propagation across turns. At each turn $s$, given a pruning budget $b$, we retain only the $b$ trajectories conditioned on the prompt whose rewards contribute most to the total reward variance within that group. The remaining trajectories, along with all their descendants, are discarded, and the process proceeds recursively until the root is reached. This approach is inspired by Xu et al. (2025a), who demonstrate that retaining trajectories with the highest reward variance within a group can reduce optimization cost while maintaining performance comparable to GRPO. We extend this principle to the multi-turn setting of Murphy.
@@ 5 Experiments
In this section, we first provide an overview of the models and datasets used in our experiments, followed by a detailed description of the setup and results, including ablation studies. We provide implementation details in App. D.
Models. We evaluate our framework using two open-source model families: Qwen3 (1.7B, 4B) (Yang et al., 2025) and OLMo2 (OLMo-2-1124-7B-Instruct) (OLMo et al., 2025). These models differ in both architecture and scale, enabling us to assess the generality of our method across diverse settings.
Datasets, Metrics, and Evaluation. We detail the datasets, metrics, and evaluation procedures employed in our experiments below.
Training Dataset: We fine-tune all models on 1,000 samples randomly drawn from the KodCode dataset (Xu et al., 2025b). This dataset was chosen due to its minimal overlap with evaluation benchmarks [^1] Refer to Section 3.2 of KodCode (Xu et al., 2025b) for details on contamination analysis.. We further study the impact of training data size in Appendix D.4.2.
Evaluation Datasets. We evaluate trained models on a suite of programming and reasoning benchmarks, including HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and BigCodeBench-Hard (Zhuo et al., 2025). For BigCodeBench-Hard, which does not include visible unit tests, we randomly sample two test cases from the full test suite to construct visible tests.
Metrics and Evaluation Protocol. Reflexion Shinn et al. (2023) is a widely adopted inference-time iterative framework that enhances reasoning by refining incorrect outputs through feedback and self-generated reflections (see App. C for details). To assess reasoning refinement and self-correction, we integrate all models into the Reflexion framework and report pass@1 under two settings: (i) Single iteration, equivalent to standard input–output prompting, and (ii) Three iterations, where feedback from visible test cases is incorporated into subsequent prompts. The agent terminates once all visible tests pass or when the maximum iteration limit is reached. Final solutions are evaluated on hidden test cases, and the resulting pass@1 is reported. Each experiment is repeated three times, and we report the mean and standard deviation of pass@1 across runs.
@@ 5.1 Murphy Experiments
We evaluate three models, Qwen3 (1.7B, 4B) (Yang et al., 2025) and OLMo-2-1124-7B-Instruct (OLMo et al., 2025), each fine-tuned on 1,000 samples randomly drawn from the KodCode dataset (Xu et al., 2025b). Evaluation is conducted on the benchmark datasets described in Sec. 5 using the Reflexion framework. We report pass@1 performance under both single- and multi-iteration settings, as summarized in Tab. 1. Additional results comparing with a naive multi-turn extension are provided in Appendix D.4.1, and experiments with higher multi-turn training are included in Appendix D.4.3. For all our tables, best results are bold and second best are underlined. Results are reported as pass@1 accuracy (% mean $\pm$ stdev).
@@
Reflexion: Multi-iteration setting. (Iter-3 in Tab. 1) We repeat the experiments with three iterations in the Reflexion framework to assess self-correction and reasoning-refinement capabilities. Increasing the number of iterations leads to consistent performance improvements across all models and benchmarks. Models trained with Murphy surpass both GRPO-trained and base models, achieving gains of up to $8\%$ over GRPO. These results demonstrate the effectiveness of multi-turn reflective optimization in enhancing reasoning refinement and self-correction.
Model Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
@@
Table 1: Performance of Qwen3-1.7B, Qwen3-4B, and OLMo-2-1124-7B-Instruct variants across evaluation benchmarks. The Rollouts column indicates the total number of generations across both stages. Murphy achieves competitive (single-turn; iter-1) or superior results (multi-turn; iter-3), with gains of up to $\mathbf{8\%}$} over GRPO under an equivalent compute budget.
Model & Strategy Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
Qwen3-1.7B
   MaRS 72 79.67 $\pm$ 3.01 86.58 $\pm$ 1.06 44.73 $\pm$ 0.50 62.00 $\pm$ 1.91 6.98 $\pm$ 3.72 20.25 $\pm$ 3.07
   MeRS ($\gamma=0.9$) 72 78.66 $\pm$ 2.11 84.76 $\pm$ 1.22 46.53 $\pm$ 0.81 60.53 $\pm$ 1.79 10.81 $\pm$ 0.00 22.52 $\pm$ 5.67
   MeRS ($\gamma=1$) 72 78.46 $\pm$ 0.70 85.57 $\pm$ 0.35 45.07 $\pm$ 1.10 60.93 $\pm$ 1.29 9.46 $\pm$ 2.44 20.27 $\pm$ 1.35
OLMo-2-1124-7B-Instruct
   MaRS 72 45.53 $\pm$ 0.70 52.24 $\pm$ 1.96 29.33 $\pm$ 0.90 39.67 $\pm$ 1.29 1.80 $\pm$ 0.39 3.38 $\pm$ 1.17
   MeRS ($\gamma=1)$ 72 37.40 $\pm$ 0.35 41.87 $\pm$ 1.27 27.87 $\pm$ 0.12 34.40 $\pm$ 2.40 0.90 $\pm$ 0.39 1.13 $\pm$ 0.39
Table 2: Ablation study comparing Max Reward (MaRS) vs Mean Reward (MeRS) propagation strategies across Qwen3-1.7B and OLMo-2-1124-7B-Instruct. MaRS consistently outperforms MeRS in both single (Iter-1) and multi-turn (Iter-3).
@@
In Subsec. 4.1, we introduced two pruning strategies: Intra-Group (IntraP) and Inter-Group (InterP) pruning. Both variants prune the rollouts tree, thereby reducing the total number of gradient updates (denoted as Updates in Tab. 3). Here, we compare the effectiveness of these strategies and observe that InterP consistently matches or surpasses the performance of Murphy without pruning. This indicates that pruning across groups offers a more robust mechanism for eliminating redundant rollouts while retaining high-quality candidates. The ablation study is conducted using the Qwen3-1.7B model, and the results are reported in Tab. 3.
Model & Strategy Updates HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
Murphy (MaRS) 72 79.67 $\pm$ 3.01 86.58 $\pm$ 1.06 44.73 $\pm$ 0.50 62.00 $\pm$ 1.91 6.98 $\pm$ 3.72 20.25 $\pm$ 3.07
Murphy (MaRS) – IntraP 36 82.34 $\pm$ 1.03 84.28 $\pm$ 2.01 45.73 $\pm$ 0.50 59.87 $\pm$ 1.70 11.03 $\pm$ 2.81 21.17 $\pm$ 2.07
Murphy (MaRS) – InterP 40 77.43 $\pm$ 2.20 86.17 $\pm$ 0.70 44.53 $\pm$ 0.90 61.20 $\pm$ 1.39 10.58 $\pm$ 2.56 24.54 $\pm$ 2.82
Table 3: Comparison of Murphy and its pruned variants on evaluation benchmarks using Qwen3-1.7B. All variants generate 72 rollouts per query. The Updates column denotes the total number of gradient steps per query. Pruned variants (IntraP, InterP) maintain competitive performance relative to the non-pruned version with significant computation reduction.
6 Conclusion
We introduced Murphy, a multi-turn reflective reinforcement learning framework that extends RLVR algorithms by incorporating iterative self-correction through both quantitative and qualitative feedback. By grounding optimization in intermediate feedback signals and propagating rewards across refinement turns, Murphy consistently enhances reasoning and code generation performance over GRPO, particularly in multi-iteration settings where feedback-driven refinement is crucial. These findings underscore the value of integrating structured feedback directly into the optimization process. Future directions include developing adaptive rollout selection strategies, extending Murphy to broader reasoning domains, and coupling it with inference-time search to further strengthen both training- and test-time reasoning.
7 Limitations
While effective, Murphy’s multi-turn design increases computational cost; pruning partially mitigates this but it remains more resource-intensive than single-stage baselines. Our experiments are limited to structured feedback in code generation, leaving open questions about generalization to noisier feedback, deeper refinement chains, and broader notions of agentic performance such as robustness and alignment with developer intent.
8 Acknowledgments
We acknowledge that Fig. 2 incorporates icons sourced from Flaticon. Full creator attributions are provided in App. G.
@@ Appendix A Extended Related Work
Appendix B GRPO: Objective and Additional Details
@@
We denote the model policy by $\pi_{\theta}(\cdot\mid\cdot)$ and the reference (older) policy by $\pi_{\theta_{\text{old}}}(\cdot\mid\cdot)$. Let $G$ be the number of generations per prompt, $\mathcal{P}(Q)$ the distribution over input prompts/questions $Q$, and $O$ the output space. For a given prompt $q\sim\mathcal{P}(Q)$, the reference policy produces a set of $G$ responses, forming a response group $\{o_{q,1},o_{q,2},\ldots,o_{q,G}\}$. Each generation $o_{q,i}\in O$ corresponds to a full output trajectory, where $o_{q,i,t}$ denotes the $t$-th token and $o_{q,i,<t}$ the prefix up to (but excluding) token $t$. We write $|o_{q,i}|$ for the sequence length of the $i$-th generation. For a given prompt $q$, the reward model assigns a scalar score to each response in the group, yielding $\mathbf{r}_{q}=\{r_{q,1},r_{q,2},\ldots,r_{q,G}\}$. Moreover, for prompt $q$, the advantage associated with the $t$-th token of the $i$-th generation is defined as $\hat{A}_{q,i,t}=(r_{q,i}-\mu(\mathbf{r}_{q}))/\sigma(\mathbf{r}_{q})$, where $\mu(\mathbf{r}_{q})$ and $\sigma(\mathbf{r}_{q})$ denote the mean and standard deviation of the group rewards, respectively. Finally, $D_{\mathrm{KL}}(\pi_{\theta}||\pi_{\theta_{\text{old}}})$ denotes the KL divergence between the current and reference policies, computed over all tokens in the generated sequences. The GRPO training objective is presented in App. B.
@@
Appendix C Reflexion

### v2

@@ 4.1 Pruning Strategies in Murphy
Intra-Group Pruning (IntraP):In IntraP, pruning operates recursively: starting from the children of level S1S-1, each pruning step is followed by reward propagation as described in Sec. 4, and the process continues backward until reaching the root. This ordering, pruning before reward reassignment, ensures that only informative trajectories are retained for credit propagation across turns. At each turn $s$, given a pruning budget $b$, we retain only the $b$ trajectories conditioned on the prompt whose rewards contribute most to the total reward variance within that group. The remaining trajectories, along with all their descendants, are discarded, and the process proceeds recursively until the root is reached. This approach is inspired by Xu et al. (2025a); they demonstrate that retaining trajectories with the highest reward variance within a group can reduce optimization cost while maintaining performance comparable to GRPO. We extend this principle to the multi-turn setting of Murphy.
@@
Table 1: Performance of Qwen3-1.7B, Qwen3-4B, and OLMo-2-1124-7B-Instruct variants across evaluation benchmarks. The Rollouts column indicates the total number of generations across two turns. Murphy (Ours) is highlighted. $\Delta_{3}$ denotes the difference between the Iter-3 performance of the GRPO/Murphy-trained models and that of Base (Iter-3), within each model block; green indicates improvement (darker = larger gain), red indicates regression.
Model Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$
Qwen3-1.7B
[…]
Qwen3-4B
[…]
OLMo-2-1124-7B-Instruct
[…]
@@ 5 Experiments
In this section, we first provide an overview of the models and datasets used in our experiments, followed by a detailed description of the setup and results, including ablation studies. We provide implementation details in App. G.
Training Dataset: We fine-tune all models on 1,000 samples randomly drawn from the KodCode dataset (Xu et al., 2025b). This dataset was chosen due to its minimal overlap with evaluation benchmarks [^1] Refer to Section 3.2 of KodCode (Xu et al., 2025b) for details on contamination analysis..
Evaluation Datasets. We evaluate trained models on a suite of programming benchmarks including HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), BigCodeBench-Hard (Zhuo et al., 2025) and Aider Polyglot [^2] https://epoch.ai/benchmarks/aider-polyglot. For BigCodeBench-Hard, which does not include visible unit tests, we randomly sample two test cases from the full test suite to construct visible tests and keep the test suite intact.
Metrics and Evaluation Protocol. Reflexion (Shinn et al., 2023) is a widely adopted inference-time iterative framework that enhances reasoning by refining incorrect outputs through iterative feedback and self-generated reflections (see App. E for details). To assess reasoning refinement and self-correction, we integrate all models into the Reflexion framework and report pass@1 under two settings: (i) Single iteration, equivalent to standard input-output prompting, and (ii) Three iterations, where feedback from evaluating visible test cases is incorporated into subsequent prompts. The agent terminates once all visible tests pass or when the maximum iteration limit is reached. Final solutions are evaluated on hidden test cases, and the resulting pass@1 is reported. Each experiment is repeated three times, and we report the mean and standard deviation of pass@1 across runs.
@@ 5.1 Murphy Experiments
We evaluate three models, Qwen3 (1.7B, 4B) (Yang et al., 2025) and OLMo-2-1124-7B-Instruct (OLMo et al., 2025), each fine-tuned on 1,000 samples randomly drawn from the KodCode dataset (Xu et al., 2025b). Evaluation is conducted on the benchmark datasets described in Sec. 5 using the Reflexion framework. We report pass@1 performance under both single and multi-iteration settings, as summarized in Tab. 1. For all our tables, best results are bold and second best are underlined. Results are reported as pass@1 accuracy (% mean $\pm$ stdev).
@@
Reflexion: Multi-iteration setting. (Iter-3 in Tab. 1) We repeat the experiments with three iterations in the Reflexion framework to assess self-correction and reasoning-refinement capabilities. Increasing the number of iterations leads to consistent performance improvements across all models and benchmarks. Models trained with Murphy surpass both GRPO-trained and base models, achieving absolute gains of up to $8\%$ over GRPO. These results demonstrate the effectiveness of multi-turn reflective optimization in enhancing reasoning refinement and self-correction.
Table 2: Ablation comparing Max Reward (MaRS) vs Mean Reward (MeRS) propagation. MaRS (Ours) is highlighted. $\Delta_{3}$ compares Iter-3 to MeRS ($\gamma=1$) Iter-3 within each model block (green indicates improvement; neutral indicates no gain).
Model & Strategy Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$
@@
[…]
Table 3: Comparison of Murphy and its pruned variants on evaluation benchmarks using Qwen3-1.7B. All variants generate 72 rollouts per query. The Updates column denotes the total number of gradient steps per query. $\Delta_{3}$ compares Iter-3 against the unpruned Murphy (MaRS) Iter-3; green indicates improvement (darker = larger gain), red indicates regression.
Model & Strategy Updates HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$
[…]
Comparisons to execution-feedback RL methods.
[…]
Generalization to Multi-Turn Code Editing.
We evaluate generalization on the Aider Polyglot Benchmark [^3] https://github.com/Aider-AI/aider/tree/main/benchmark, which tests self-correction through execution feedback: models receive two attempts per problem, with unit test results provided as feedback after a failed first attempt. This setup closely mirrors Murphy’s training objective. Without additional training, we evaluate the same Qwen3-4B checkpoints from our main experiments (Tab. 1) following the official evaluation setup. The Murphy-trained model achieves 10.9% pass@1 and 21.6% pass@2, outperforming both the base (7.6% / 18.7%) and GRPO-trained (8.0% / 17.8%) models, confirming that Murphy’s gains transfer to external feedback-driven tasks.
Additional Experiments. Further analysis is provided in the appendices: effect of training data size (App.G.4.2), computational cost (App.F), comparison with naive multi-turn GRPO (App.G.4.1), and training beyond two turns (App.G.4.3). App. D confirms that Murphy’s gains generalize across different iterative scaffolds.
@@
In Subsec. 4.1, we introduced two pruning strategies: Intra-Group (IntraP) and Inter-Group (InterP) pruning. Both reduce the rollout tree size, thereby decreasing the total number of gradient updates (Tab. 3). We compare their effectiveness using Qwen3-1.7B. While both strategies reduce computational cost, they exhibit different trade-offs. IntraP improves single-turn performance (Iter-1) across all benchmarks but shows modest regression on multi-turn evaluation (Iter-3) for HumanEval and MBPP. In contrast, InterP maintains comparable multi-turn performance on HumanEval and MBPP while achieving notable gains on BigCodeBench. These results suggest that InterP offers a more robust trade-off, preserving multi-turn self-correction capabilities while reducing computational cost.
6 Conclusion & Limitations
We introduced Murphy, a multi-turn reflective reinforcement learning framework that extends RLVR algorithms by incorporating iterative self-correction through both quantitative and qualitative feedback. By grounding optimization in intermediate feedback signals and propagating rewards across refinement turns, Murphy consistently improves reasoning and code generation performance, particularly in multi-iteration settings where feedback-driven refinement is crucial. These findings underscore the value of integrating structured feedback directly into the optimization process. While effective, Murphy’s multi-turn design increases computational cost; pruning partially mitigates this but it remains more resource-intensive than single-turn baselines. We study structured execution feedback for code, and generalization to noisier feedback, deeper refinement, and broader agentic objectives is open. In future work, we aim to develop adaptive turn/rollout selection and tool-augmented optimization (retrieval, execution, APIs).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ 6 Conclusion & Limitations
This appendix complements the main text with extended related work, formal definitions, and additional analyses and ablations. The contents are organized as follows:
  • App. A: Extended discussion of related work, covering LLM agents for software development and reinforcement learning with verifiable rewards.
  • App. B: Detailed treatment of the multi-turn rollout tree formalism and indexing conventions.
  • App. C: Review of the GRPO objective and associated notation.
  • App. D: Sensitivity analysis of Murphy-trained models to different multi-iteration inference scaffolds.
  • App. E: Description of the Reflexion framework used for evaluation.
  • App. F: Computational cost comparison between multi-turn and single-turn training.
  • App. G: Implementation details, hyperparameters, and additional experiments including ablations on dataset size, naive multi-turn baselines, and pruning strategies.
  • App. H: Prompts used in our experiments
  • App. I: Potential risks
  • App. J: Icon attributions for figures in the main text.
@@ Appendix A Extended Related Work
Appendix B Multi-turn Rollout Formalism (Detailed)
This appendix section provides a detailed treatment of the rollout-tree construction and indexing used throughout the main paper.
We define a feedback-conditioned rollout tree that captures how model generations evolve across multiple turns of interaction with the environment. Let $s$ denote the turn index, $S$ the total number of turns, and $G_{s}$ the number of generations per prompt at turn $s$.
Turn 1:In the first turn ($s=1$), the model receives an input prompt $q_{(1)}$ sampled from $\mathcal{P}(Q)$ and generates $G_{1}$ candidate programs which forms a response group $\{o_{\{q_{(1)},1\}},o_{\{q_{(1)},2\}},\ldots,o_{\{q_{(1)},G_{1}\}}\}$. Note that $\{o_{\{q_{(1)},j\}}\}_{j=1}^{G_{1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(1)})$. Each generation is executed against its associated test suite, producing a numerical reward $r_{\{q_{(1)},j\}}$ and environment feedback $f_{\{q_{(1)},j\}}$. The reward represents the proportion of test cases passed. The environment feedback contains the specific unit tests that passed or failed, along with any corresponding error messages. These $G_{1}$ generations form the first layer of output nodes in the rollout tree.
Turn $s$:Building on the previous turn, this procedure extends recursively to any turn $s\in\{1,\dots,S-1\}$. We define $i_{[1:s]}=i_{1},\dots,i_{s}$ as a sequence of branch indices that trace a specific path through the tree, where $i_{j}$ indicates the $i$’th candidate that was selected at turn $j$. Similarly, $i_{[1:1]}$ denotes $i_{1}$. For each generation $o_{{q_{(s,i_{[1:s-1]})},i_{s}}}$ that fails to achieve the maximum reward, we construct a feedback-conditioned prompt:
$q_{(s+1,i_{[1:s]})}=[q_{(s,i_{[1:s-1]})},\,o_{{q_{(s,i_{[1:s-1]})},i_{s}}},\,f_{{q_{(s,i_{[1:s-1]})},i_{s}}}]$
where $\cdot$] denotes textual concatenation. The model is then re-invoked to generate $G_{s+1}$ new candidate solutions:
$\{o_{\{q_{(s+1,i_{[1:s]})},k\}}\}_{k=1}^{G_{s+1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(s+1,i_{[1:s]})})$
Each candidate is evaluated to obtain its reward and feedback $(r_{\{q_{(s+1,i_{[1:s]})},k\}},\,f_{\{q_{(s+1,i_{[1:s]})},k\}})$. The resulting generations, $\{o_{\{q_{(s+1,i_{[1:s]})},k\}}\}_{k=1}^{G_{s+1}}$, form the child nodes of parent node $o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$.
A complete path from the root (the initial prompt at turn 1) to a leaf at turn $S$ (final output) can be expressed as:
$q_{(1)}\rightarrow o_{{q_{(1)},i_{1}}}\rightarrow q_{(2,i_{[1:1]})}\rightarrow o_{{q_{(2,i_{[1:1]})},i_{2}}}\rightarrow\cdots\rightarrow o_{\{q_{(S,i_{[1:S-1]})},i_{S}\}}$
where the indices $i_{1},\dots,i_{S}$ specify which branch is taken at each turn. Leaf nodes at turn $S$ represent the final generations obtained after completing all refinement steps. Once rewards for all turns are computed, the rewards from these terminal nodes are propagated backward through their ancestors according to the credit assignment strategies described below.
Appendix C GRPO: Objective and Additional Details
@@
We denote the model policy by $\pi_{\theta}(\cdot\mid\cdot)$ and the reference (older) policy by $\pi_{\theta_{\text{old}}}(\cdot\mid\cdot)$. Let $G$ be the number of generations per prompt, $\mathcal{P}(Q)$ the distribution over input prompts/questions $Q$, and $O$ the output space. For a given prompt $q\sim\mathcal{P}(Q)$, the reference policy produces a set of $G$ responses, forming a response group $\{o_{q,1},o_{q,2},\ldots,o_{q,G}\}$. Each generation $o_{q,i}\in O$ corresponds to a full output trajectory, where $o_{q,i,t}$ denotes the $t$-th token and $o_{q,i,<t}$ the prefix up to (but excluding) token $t$. We write $|o_{q,i}|$ for the sequence length of the $i$-th generation. For a given prompt $q$, the reward model assigns a scalar score to each response in the group, yielding $\mathbf{r}_{q}=\{r_{q,1},r_{q,2},\ldots,r_{q,G}\}$. Moreover, for prompt $q$, the advantage associated with the $t$-th token of the $i$-th generation is defined as $\hat{A}_{q,i,t}=(r_{q,i}-\mu(\mathbf{r}_{q}))/\sigma(\mathbf{r}_{q})$, where $\mu(\mathbf{r}_{q})$ and $\sigma(\mathbf{r}_{q})$ denote the mean and standard deviation of the group rewards, respectively. Finally, $D_{\mathrm{KL}}(\pi_{\theta}||\pi_{\theta_{\text{old}}})$ denotes the KL divergence between the current and reference policies, computed over all tokens in the generated sequences. The GRPO training objective is presented in App. C.
@@
Appendix D Sensitivity to Multi-Iteration Scaffolds
Murphy is explicitly designed to train models to incorporate environment feedback and self-correct over multiple turns. In a single-iteration evaluation setting, where no feedback is available, Murphy-trained models naturally perform comparably to GRPO-trained models. In contrast, under multi-iteration evaluation—typical of agentic settings where feedback is present—Murphy yields substantially larger gains. This behavior is expected and reflects the objective of Murphy, which is to improve a model’s ability to utilize feedback across turns rather than to optimize single-pass generation.
To assess sensitivity to the choice of multi-iteration scaffold, we conduct additional experiments using the Qwen3-1.7B model. We compare the base model, the GRPO-trained model, and the Murphy-trained model under three distinct scaffolds: Reflexion (Shinn et al., 2023), LATS (Zhou et al., 2024), and DoT (Lingam et al., 2025). Importantly, no additional training is performed. We evaluate the same GRPO checkpoint reported in Table 1 and the same Murphy (MaRS, IntraP) checkpoint reported in Table 3. All scaffolds use their standard configurations with three iterations. We report mean accuracy and standard deviation over three independent runs on HumanEval.
Table 4: Sensitivity of Qwen3-1.7B performance on HumanEval to different multi-iteration scaffolds. No additional training is performed.
Model Reflexion LATS DoT
Qwen3-1.7B (Base) 80.07 $\pm$ 1.27 78.95 $\pm$ 1.29 78.35 $\pm$ 3.88
Qwen3-1.7B GRPO 82.11 $\pm$ 1.76 80.79 $\pm$ 3.01 83.74 $\pm$ 0.57
Qwen3-1.7B Murphy (MaRS) 84.28 $\pm$ 2.01 86.89 $\pm$ 0.42 85.66 $\pm$ 0.43
Across all scaffolds, the Murphy-trained model consistently outperforms both the GRPO-trained model (trained under a matched compute budget) and the base model. These results indicate that the gains from Murphy are not dependent on any particular inference-time scaffold, but instead reflect improved feedback utilization learned during training.
Appendix E Reflexion